26millecai:

The leg of a right triangle is 3 units and the hypotenuse is 6 units. What is the length, in units, of the other leg of the triangle? (5 points) 3 units 9 units Square root of 18 units Square root of 27 units

QuestionCoveBot:

Since the hypotenuse is 6 units, and the one of the legs is 3, what would guess the other leg unit is?

26millecai:

3?

QuestionCoveBot:

Correct. Both legs are 3 units each, making the hypotenuse 6 units.

26millecai:

so thats the answer?

QuestionCoveBot:

Yes

26millecai:

Thank you!!!

QuestionCoveBot:

You're welcome.

snowflake0531:

Pythagorean theorem: \(a^2+b^2=c^2\) where a and b are hte legs and c is the hypotenuse replace variables: \(3^2 + b^2 = 6^2 \\9+b^2=36\\b^2=27\\b=\sqrt{27} \)